OpenStudy (anonymous):

Find A when P=4000 r =45 n=12 t=10 A=P(1+r/n)^nt

OpenStudy (anonymous):

r was supposed to equal 4% that was a typo

jimthompson5910 (jim_thompson5910):

A=P(1+r/n)^(nt) A=4000(1+0.04/12)^(12*10) A=4000(1+0.00333333333333)^(120) A=4000(1.00333333333333)^(120) A=4000(1.49083268241762) A=5963.33072967048 A=5963.33